OpenStudy (rowanwhitethorn):

help solving a compound inequality 4x≤ 12 and -7x≤ 21

satellite73 (satellite73):

first one \[4x\leq 12\] divide both sides by \(4\)

satellite73 (satellite73):

second one \[-7x\leq 21\] divide both sides by \(-7\) and change the inequality from \(\leq\) to \(\geq\)

satellite73 (satellite73):

let me know what you get